# GFlowGR: Fine-tuning Generative Recommendation Frameworks with Generative Flow Networks

### v1

@@ Abstract.
Generative recommendations (GR), which usually include item tokenizers and generative Large Language Models (LLMs), have demonstrated remarkable success across a wide range of scenarios. The majority of existing research efforts primarily concentrate on developing powerful item tokenizers or advancing LLM decoding strategies to attain superior performance. However, the critical fine-tuning step in GR frameworks, which is essential for adapting LLMs to recommendation data, remains largely unexplored. Current approaches predominantly rely on either the next-token prediction loss of supervised fine-tuning (SFT) or recommendation-specific direct preference optimization (DPO) strategies. Both methods ignore the exploration of possible positive unobserved samples, which is commonly referred to as the exposure bias problem. To mitigate this problem, this paper treats the GR as a multi-step generation task and constructs a GFlowNets-based fine-tuning framework (GFlowGR). The proposed framework integrates collaborative knowledge from traditional recommender systems to create an adaptive trajectory sampler and a comprehensive reward model. Leveraging the diverse generation property of GFlowNets, along with sampling and heuristic weighting techniques, GFlowGR emerges as a promising approach to mitigate the exposure bias problem. Extensive empirical results on two real-world datasets and with two different GR backbones highlight the effectiveness and robustness of GFlowGR.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large Language Models (LLMs) possess exceptional capabilities in understanding, reasoning, and generation. As a result, they are increasingly utilized across various domains, including healthcare, education, and science (Naveed et al., 2023), as well as for information retrieval and recommendations in numerous online services, such as E-commerce (Li et al., 2024a; Zhai, 2024). Notably, the impressive generative capabilities of LLMs have spurred the development of generative recommendation (GR) frameworks (Rajput et al., 2023; Huang et al., 2024; Hua et al., 2023). As visualized in Figure 1 (a), these frameworks typically consist of three key components (Li et al., 2024b): item tokenizers, user prompts, and generative LLMs. Meanwhile, the training paradigms usually consist of two stages: optimizing the item tokenizer to represent items with several tokens and fine-tuning the LLMs for recommendation objectives like next-item prediction.
Existing research on GR has primarily concentrated on enhancing the item tokenizer (Zheng et al., 2025; Liu et al., 2024c; Lin et al., 2025b; Wang et al., 2024a; Zhu et al., 2024) and the associated optimization processes, designing user prompts (Liu et al., 2024a; Wang et al., 2024b), and advancing generative LLMs (Lin et al., 2025c; Ding et al., 2024). However, there has been little focus on innovating the fine-tuning step of LLMs. While this step is crucial for GR frameworks: the semantics of recommendations can significantly differ from those of LLMs pretrained for general language purposes (Lin et al., 2025a; Wu et al., 2024). For instance, terms like “Comment” represent specific user behaviors within recommendation data, even though they are commonly used verbs in general language tasks. Moreover, the item identifier in GR may consist of out-of-vocabulary tokens for pretrained LLMs (Huang et al., 2024; Wang et al., 2024a; Rajput et al., 2023), which are unlikely to be generated by zero-shot LLMs.
Current LLM fine-tuning techniques applied for GR can be generally divided into two categories (Wu et al., 2025; Wang et al., [n. d.]): supervised fine-tuning (SFT) and reinforcement learning-based fine-tuning (RLFT). The former optimizes LLMs with the next token prediction loss (Brown et al., 2020; Radford et al., 2019; Hu et al., 2022) and handcrafted auxiliary alignment loss (Liu et al., 2024a; Zheng et al., 2024; Wang et al., 2024b), which focuses solely on positive samples within the dataset. On the one hand, SFT does not incorporate any negative samples for comparison with the positive ones, leaving a gap with the ranking objective of recommender systems (Chen et al., [n. d.]). On the other hand, SFT cannot explore potential positive samples to extend the collected data, which is always sparse (Gao et al., 2022). With the exploration ability, RLFT methods like proximal policy optimization (PPO) (Schulman et al., 2017), and direct preference optimization (DPO) (Rafailov et al., 2023) can improve the model capacity on the basis of SFT by achieving the maximum reward. For example, recent DPO-based methods facilitate LLM fine-tuning by integrating augmented negative samples for GR (Chen et al., [n. d.]; Bai et al., 2024; Liao et al., 2024; Gao et al., 2025a; Deng et al., 2025), enabling LLMs to rank candidate items. However, both solutions cannot match the diverse nature of recommender systems that users may prefer multiple items, while only one among them can be observed. SFT regards the unobserved interactions as negative, and RLFT methods predefine the largest reward to the observed samples. This reliance on the collected data that fails to fully reveal the real user preference can lead to biased user preferences, known as exposure bias (Chen et al., 2023). Therefore, it calls for an LLM fine-tuning solution that can explore potential positive samples and leverage augmented samples to achieve diverse results.
Recent advancements in Generative Flow Networks (GFlowNets) present a promising solution: by training LLMs to generate token sequences with probabilities proportional to a specific reward distribution (Bengio et al., 2023), GFlowNets can identify potential positive candidates and thus extend the recommendation diversity. This capability is supported by recent studies demonstrating the application of GFlowNets for multi-step reasoning tasks (Kang et al., 2025; Takase et al., 2024; Yu et al., 2024) and compositional generation (Jain et al., 2022). Generative recommendation tasks are inherently similar to these tasks; as illustrated in Figure 1 (b), recommending an item involves multiple generation steps of a single token.However, applying GFlowNets for GR still faces significant challenges: how to organize training trajectories (samples) and establish a reward model to assess the terminal state for GFlowNets. Specifically, the collected data consists solely of positive interaction logs from users, resulting in sparse training trajectories for GFlowNets. This scenario necessitates high-quality augmented trajectories for robust training. Furthermore, merely treating these augmented trajectories as negative samples and assigning them low rewards is inadequate for overcoming the exposure bias problem, as these augmented trajectories are not consistently negative.
To address these limitations, this paper constructs the GFlowNets-based fine-tuning paradigm for GR (GFlowGR). Drawing inspiration from prior insights on aligning LLMs with collaborative recommendation models (Hong et al., [n. d.]; Yang et al., 2024; Lin et al., 2025a), we incorporate collaborative scores when designing the trajectory sampling and reward model for GFlowGR. Specifically, we develop an adaptive trajectory sampler that generates easy-to-hard augmented trajectories, thereby implementing a curriculum-style training approach (Wang et al., 2021a) to consistently provide high-quality training data. Additionally, we construct a multi-faceted reward system that thoroughly considers augmentation signals, collaborative scores, and semantic similarities to effectively differentiate possibly preferred items from the sampled unobserved trajectories. Finally, to further mitigate the exposure bias problem, we assign confidence weights to each sample based on the agreement between augmentation signals and collaborative scores. We summarize the major contributions of this paper as follows:
  • We formulate GR as a multi-step generation process and develop a GFlowNets-based fine-tuning framework, GFlowGR. This framework helps mitigate the exposure bias problem by learning generation probabilities that are proportional to the reward.
  • We implement an adaptive trajectory sampler to select high-quality training samples for GFlowGR and establish a comprehensive reward model to consider various reward signals effectively.
  • Our extensive experimental results, obtained from two public benchmarks using two backbone GR models, demonstrate the effectiveness and versatility of GFlowGR.
Figure 1. Overview of GR and GFlowNets, with the length of item identifiers $L=3$.
@@ 2. Preliminary
Generative Recommendation (GR). In this paper, we define GR frameworks by referencing recent works (Rajput et al., 2023; Wang et al., 2024a). A typical GR workflow is illustrated in Figure 1 (a). The framework first constructs an item tokenizer using item features to decompose items $v$ into token sequences $\{t_{l}\}_{l<L}$, serving as item identifiers. The lower portion of the figure exemplifies that a skyblue shirt is encoded as the token sequence $<a\_1><b\_4><c\_2>$ based on its title and descriptions. To infer the user preference, GR frameworks assemble user prompts (LLM inputs) by integrating task descriptions, historical user interactions, and profile information (e.g., age, gender denoted as $<\underline{X}>$). These prompts, exemplified by tokenized interaction histories (blue tokens for jackets, shirts, pants) and profiles, are fed into LLMs to generate token sequences representing the most probable next item. As shown in the upper portion of Figure 1 (a), this process produces $<a\_2><b\_4><c\_10>$, decoding to a blue polo shirt as the recommended item.
To achieve the recommendation goal, the training paradigm of GR comprises two stages: optimizing the item tokenizer ($\mathcal{L}_{T}$) and the generative LLMs ($\mathcal{L}_{\text{GR}}$), respectively. For the tokenizer, one prevalent tokenization strategy is based on RQ-VAE solutions (Lee et al., 2022; Rajput et al., 2023), where $\mathcal{L}_{T}$ includes a codebook loss and a reconstruction loss. Studies enhancing the tokenizer often incorporate auxiliary losses into $\mathcal{L}_{T}$, such as diversity (Wang et al., 2024a) and collaborative-constrained losses (Lin et al., 2025b). $\mathcal{L}_{\text{GR}}$ typically adopts the next-token prediction loss (Deng et al., 2025; Rajput et al., 2023; Wang et al., 2024a; Liu et al., 2024a)::::
@@
Motivation. Two key insights motivate us to apply GFlowNets for GR: first, GR inherently involves a multi-step generation process well-suited for GFlowNet modeling. As shown in Figure 1 (b), generating the next item identifier unfolds over $L$ sequential steps: starting from an initial state $s_{0}=\{\emptyset\}$, LLMs iteratively generate tokens (e.g., $<a\_2>$, $<b\_4>$, $<c\_10>$) to form intermediate states $s_{1}=\{<a\_2>\}$, $s_{2}=\{<a\_2><b\_4>\}$, and finally reach the terminal state $s_{3}=\{<a\_2><b\_4><c\_10>\}$, which decodes to a specific item (e.g., a blue polo). This sequential structure aligns naturally with GFlowNets’ trajectory-based framework. Second, GFlowNets offer two critical advantages for GR: (a) Diverse High-Reward Generation: by linking generation probability to reward ($\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$), GFlowNets can explore and generate diverse items with high predicted rewards. Theoretical analyses (Bengio et al., 2021, 2023) and applications in domains like DNA design (Jain et al., 2022) and multi-step reasoning (Kang et al., 2025; Yu et al., 2024) validate this capability. In recommendation, this addresses exposure bias by treating unobserved interactions as potential positives rather than default negatives. Particularly, during training, GFlowNets can identify augmented items with high reward scores as potential positive and prioritize them over low-reward alternatives. (b) Off-Policy Data Utilization: GFlowNets enable off-policy learning, allowing GR frameworks to incorporate augmented interactions into training. This extends SFT, which relies solely on collected positive samples, to leverage a broader range of potential user preferences. The lower part of Figure 1 (b) illustrates these: by associating generation probabilities with rewards, GFlowNets empower GR to prioritize the unseen blue jacket (aligned with the user’s formal dressing history) over the yellow undershirt, even when both are unobserved. This contrasts with conventional methods that treat unobserved items equally, thereby enhancing diversity and mitigating exposure bias.
@@ 3. Methodology
In this section, we detail the proposed GFlowGR framework, including its adaptive trajectory sampler, comprehensive reward model, and training/inference procedures. The advantages of GFlowGR in addressing exposure bias and negative sampling are summarized in the Supplementary Material.
@@
The architecture of GFlowGR is depicted within the dashed block of Figure 2. Notably, GFlowGR modifies only the LLM fine-tuning stage for GR while keeping the inference pipeline unchanged. The framework integrates a collaborative model (CM) that is capable of scoring candidate items (blue dotted block) based on user interaction histories (pink dotted block) as a core component of both the trajectory sampler and reward model. The overall workflow is as follows: for a given input pair $(U,v)$, the adaptive trajectory sampler selects augmented samples $v^{\prime}$ as extended generation trajectories. GFlowNet then estimates the forward probabilities with $\mathbb{P}_{F}$ and flow values with $\mathcal{F}$ for both the original trajectory $\tau$ (corresponding to the ground-truth $v$) and augmented trajectories $\tau^{\prime}$ (corresponding to $v^{\prime}$). The reward model $\mathcal{R}$ integrates collaborative scores ($r_{c}$), augmentation signals from the trajectory sampler ($r_{a}$), and token similarities between positive and augmented item identifiers ($r_{\text{sim}}$) to compute rewards. Finally, GFlowGR computes the GFlowNet loss $\mathcal{L}_{\text{GFN}}$ for the LLM fine-tuning.
@@
Figure 2. Framework of GFlowGR. The visualized example shows a positive sample (blue polo shirt) and an augmented sample (yellow undershirt). Forward probabilities ($\mathbb{P}_{F}$) for both trajectories are displayed (e.g., $0.9,0.8,0.83$ for $\tau$ ), while estimated flow values ($\mathcal{F}$) are omitted for brevity.
3.1. Adaptive Trajectory Sampler

### v2

@@ Abstract.
Generative recommendation (GR) systems, which leverage Large Language Models (LLMs) to produce recommendations, have demonstrated remarkable success in industrial applications. However, despite significant research into item tokenization and decoding strategies, a critical gap persists between the standard training paradigm and the demands of real-world deployment. Existing methods often fail to effectively leverage the rich, value-aware signals in user interaction data. They typically fine-tune models on a single ground-truth item in a value-unaware manner, while even reward-based approaches often neglect crucial token-level supervision, ultimately limiting performance. To address these limitations, we formulate the GR task as a multi-step generation problem and introduce GFlowGR, a novel fine-tuning framework based on Generative Flow Networks (GFlowNet). The modular design of GFlowGR features three core components: a trajectory sampler, a reward model, and the GFlowNet loss. We present the various configurations of GFlowGR for diverse practical scenarios. Extensive experiments on three real-world datasets using two GR backbones demonstrate the effectiveness of GFlowGR. For real-world deployment, GFlowGR is live across all of Taobao’s advertising businesses, driving a 1% relative increase in billion-level annual revenue since 2025 May.
@@
copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAEisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large Language Models (LLMs) possess exceptional capabilities in understanding, reasoning, and generation. As a result, they are increasingly utilized across various domains, including healthcare, education, and science (Naveed et al., 2023), as well as for information retrieval and recommendations in numerous online services, such as E-commerce (Li et al., 2024a; Zhai, 2024). Notably, the impressive generative capabilities of LLMs have spurred the research on generative recommendation (GR) frameworks (Rajput et al., 2023; Huang et al., 2024; Hua et al., 2023), as well as the increasing deployment in industry (Zhou et al., 2025; Zheng et al., 2025b; Guo et al., 2025; Wei et al., 2025).
As visualized in Figure 1 (a), these frameworks typically consist of three key components (Li et al., 2024b): item tokenizers, user prompts, and generative LLMs. Despite significant advances in item tokenization (Zheng et al., 2025a; Liu et al., 2024c; Lin et al., 2025b; Wang et al., 2024a; Zhu et al., 2024), user prompting, and inference optimization (Lin et al., 2025c; Ding et al., 2024), a critical bottleneck for deploying industrial-scale GR systems remains the LLM finetuning process.
The standard paradigm, Supervised Fine-Tuning (SFT) (Rajput et al., 2023), is fundamentally misaligned with the core objective of production systems. These systems, whether deployed for retrieval (Yang et al., 2025) or as end-to-end “OneModels” (Zhou et al., 2025), are tasked with generating a diverse and high-quality set of recommendations, where items may have different values (e.g., interaction type for training and recommendation scores for prediction). However, SFT’s sequence-to-sequence objective is problematic for two key reasons. First, it trains the model to predict only a single ground-truth item per instance, ignoring the rest of the relevant item set. Second, it often fails to model the varying value across interacted items. These create a significant gap between offline training and real-world deployment objectives for online serving.
To bridge this gap, recent work has turned to Reinforcement Learning (RL) to better exploit the full, value-aware item set (Chen et al., 2025; Zhou et al., 2025). For example, Chen et al. (2025) adapt Direct Preference Optimization (DPO) (Rafailov et al., 2023) by constructing preference pairs from the user’s interaction history; items with higher scores (e.g., from longer dwell times or clicks) are treated as the “chosen” examples, while those with lower scores become the “rejected” ones. Similarly, OneRec-V2 (Deng et al., 2025) modifies the group-wise learning objective GRPO (Shao et al., 2024) to leverage the entire set of real-world user interactions. A common limitation of these methods, however, is their reliance on outcome-based rewards. They assign rewards at the item level but fail to supervise the underlying token-level generation process. Since items in GR systems are composed of multiple tokens (Figure 1 (b)), this lack of process-level supervision represents a critical performance bottleneck (Gao et al., 2025b). Therefore, it calls for an effective fine-tuning solution that not only learn from a value-aware item set but also provide direct supervision over the token-by-token generation process.
Recent advancements in Generative Flow Networks (GFlowNets) present a promising solution: Organizing the item set as generation trajectories, assigning item value as the termination reward, GFlowNets can attribute flow values as the token-level reward, which also provide the additional benefit of generating token sequences with probabilities proportional to the reward distribution (Bengio et al., 2023), enhancing the generation diversity, which is also a practical demand of industrial recommendation scenarios. This capability is supported by recent studies demonstrating the application of GFlowNets for multi-step reasoning tasks (Kang et al., 2025; Takase et al., 2024; Yu et al., 2024) and compositional generation (Jain et al., 2022). Generative recommendation tasks are inherently similar to these tasks; as illustrated in Figure 1 (b), recommending an item involves multiple generation steps of a single token. Despite the recent successes in deploying GFlowNets for traditional recommendations (Liu et al., 2023, 2024b) or text-based LLM recommendations (Gao et al., 2025b), there still lack a conceptual GFlowNets framework for GR systems.
Consequently, this paper constructs the GFlowNet-based fine-tuning paradigm for GR (GFlowGR). We suggest three key components for GFlowGR: a trajectory sampler to structure training data, a reward model to estimate item value, and a GFlowNet training objective. We detail the conceptual implementation of these components and show how they assemble into a cohesive system. Designed for practical adoption, GFlowGR can cooperate with existing recommendation models, and serve as a powerful supplement to standard SFT. Since May 2025, GFlowGR has been deployed in Taobao’s production GR system, serving hundreds of millions of daily active users and driving billions of dollars in annual advertising revenue. We summarize the major contributions of this paper as follows:
  • We formulate GR as a multi-step generation problem and introduce GFlowGR, a new fine-tuning paradigm based on GFlowNets that provides token-level supervision.
  • We present the modular architecture of GFlowGR, consisting of a trajectory sampler, reward model, and GFlowNet loss. We demonstrate its flexibility by instantiating these components with various configurations.
  • We show through extensive experiments on three public datasets and two backbone models that GFlowGR consistently and significantly outperforms existing methods.
  • We detail the successful deployment of GFlowGR in Taobao’s production system, where it has been serving hundreds of millions of daily active users and driving billions of annual revenue.
Figure 1. Overview of GR and GFlowNets, with the length of item identifiers $L=3$. (a) A process for constucting the GR framwork. (b) Generative flow structure is exemplified with a set of user interaction, which typically includes positive feedback (“Clicked”), impression (“Presented”), and unexposed items (“unpresented”).
@@ 2. Preliminary
Generative Recommendation (GR). In this paper, we define GR frameworks by referencing recent works (Rajput et al., 2023; Wang et al., 2024a). A typical GR workflow is illustrated in Figure 1 (a). The framework first constructs an item tokenizer using item features to decompose items $v$ into token sequences $\{t_{l}\}_{l<L}$, serving as item identifiers. The lower portion of the figure exemplifies that a skyblue shirt is encoded as the token sequence <a_1><b_4><c_2><a\_1><b\_4><c\_2> based on its title and descriptions. To infer the user preference, GR frameworks assemble user prompts (LLM inputs) by integrating task descriptions, historical user interactions, and profile information (e.g., age, gender denoted as $\underline{X}>$). These prompts, exemplified by tokenized interaction histories (blue tokens for jackets, shirts, pants) and profiles, are fed into LLMs to generate token sequences representing the most probable next item. As shown in the upper portion of Figure 1 (a), this process produces <a_2><b_4><c_10><a\_2><b\_4><c\_10>, decoding to a blue polo shirt as the recommended item.
To achieve the recommendation goal, the training paradigm of GR comprises two stages: optimizing the item tokenizer ($\mathcal{L}_{T}$) and the generative LLMs ($\mathcal{L}_{\text{GR}}$), respectively. For the tokenizer, one prevalent tokenization strategy is based on RQ-VAE solutions (Lee et al., 2022; Rajput et al., 2023), where $\mathcal{L}_{T}$ includes a codebook loss and a reconstruction loss. Studies enhancing the tokenizer often incorporate auxiliary losses into $\mathcal{L}_{T}$, such as diversity (Wang et al., 2024a) and collaborative-constrained losses (Lin et al., 2025b). $\mathcal{L}_{\text{GR}}$ typically adopts the next-token prediction loss (Deng et al., 2025; Rajput et al., 2023; Wang et al., 2024a; Liu et al., 2024a)::
@@
Motivation. Two key insights motivate us to apply GFlowNets for GR: first, GR inherently involves a multi-step generation process well-suited for GFlowNet modeling. As shown in Figure 1 (b), generating the next item identifier unfolds over $L$ sequential steps: starting from an initial state $s_{0}=\{\emptyset\}$, LLMs iteratively generate tokens (e.g., <a_2><a\_2>, <b_4><b\_4>, <c_10><c\_10>) to form intermediate states $s_{1}=\{<a\_2>\}$, $s_{2}=\{<a\_2><b\_4>\}$, and finally reach the terminal state $s_{3}=\{<a\_2><b\_4><c\_10>\}$, which decodes to a specific item (e.g., a blue polo). This sequential structure aligns naturally with GFlowNets’ trajectory-based framework. Second, GFlowNets offer two critical advantages for GR: (a) Item Value Perception: By linking generation probability to reward ($\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$), GFlowNets can explore and generate diverse items with high rewards (i.e., values). Theoretical analyses (Bengio et al., 2021, 2023) and applications in domains like DNA design (Jain et al., 2022) and multi-step reasoning (Kang et al., 2025; Yu et al., 2024) validate this capability. This addresses the need for ranking ability in recommendations: preferred items are required to possess a higher rank than low-value items. (b) Full Set Utilization: GFlowNets enable the learning from a sort of trajectories, allowing GR frameworks to incorporate all interactions into the single training step. This extends SFT, which relies solely on the single positive sample, to leverage a broader range of user interactions. The lower part of Figure 1 (b) illustrates these: by associating generation probabilities with rewards/values and including all interactions (clicked, presented but not clicked, and unpresented), GFlowNets empower GR to prioritize the presented blue jacket (have a larger probability to be clicked) over the yellow undershirt with a lower reward for unpresented items.
@@ 3. Methodology
In this section, we detail the proposed GFlowGR framework, including the organization of trajectory sampler, reward model, and GFN losses. We also present the possible integration of existing collaborative models and elaborate the training and inference procedures.
@@
The architecture of GFlowGR is depicted within Figure 2. Notably, GFlowGR modifies only the LLM fine-tuning stage for GR while keeping the inference pipeline unchanged. The overall workflow is as follows: for a given input pair $(U,v)$, where $v$ denotes the single item with the positive feedback [^1] In fact, the number of positive items is usually larger than one in the session. We discuss with the single positive situation for clarity.. Trajectory sampler selects augmented samples or loads the collected system logs to extend the item set as $\{v_{n}\}_{n\leq N})$, where $v_{1}=v$. GFlowNet then estimates the forward probabilities with $\mathbb{P}_{F}$ and flow values with $\mathcal{F}$ for all trajectories $\tau_{n}$ (corresponding to $v_{n}$). The reward model then assigns a reward $\mathcal{R}$ for each trajectory. Finally, GFlowGR computes the GFlowNet loss $\mathcal{L}_{\text{GFN}}$ for the LLM fine-tuning.
It is also possible for the framework to integrate a existing collaborative model (CM): for the candidate item $v_{n}$, CM is capable of providing the score based on user interaction histories for both the trajectory sampler and reward model.
@@
Figure 2. Framework of GFlowGR. The visualized example shows a the learning on a NN-item set illustrated with two items among the set, including a positive sample (blue polo shirt) and an unpresented sample (yellow undershirt). Forward probabilities ($\mathbb{P}_{F}$) for both trajectories are displayed (e.g., $0.9,0.8,0.83$ for $\tau_{1}$ ), while estimated flow values ($\mathcal{F}$) are omitted for brevity. The bluesky shaded area presents a example of incorporating collaborative models (CM).
3.1. Trajectory Sampler
This section details the construction of the trajectory sampler, which produces the learning trajectories for a given user-interaction pair $(U,v)$. There are four methods for constructing the trajectory set: interaction log, random sampling, CM-based sampling, and on-policy sampling.
1. Interaction log. For the industrial system, it is straightforward to cutoff the user interaction log for the trajectory set, which typically contains more than $N$ samples. Specifically, with this solution, the trajectory sampler simply finds the session for $(U,v)$ and selects the N1N-1 samples to formulate the item set.